What is the pH when 28.00 mL of 0.200 M potassium hydroxide (KOH) have been added to 50.00 mL of 0.100 M hydrochloric acid (HCl) solution?


1
Expert's answer
2021-01-22T06:25:18-0500

"HCl + KOH \to KCl + H_2O"


number of moles of KOH = 28/1000 × 0.200 = 0.0056 moles.

number of moles of HCl = 50/1000 × 0.100 = 0.0050 moles.


Therefore, KOH is in excess and 0.0006 moles of it will be left after a complete neutralisation.

The total volume of the solution after the reaction is 28.00mL + 50.00mL = 78.00mL


"KOH \rightleftarrows K^+ + OH^-"

since 1 mole of KOH produces 1 mole of hydroxyl ion,

the number of moles of hydroxyl ion is 0.0006 moles


[OH-] = 0.0006/0.0078 = 0.077M

pOH = -log(0.077) = 1.11


pH + pOH = 14

pH = 14 - 1.11 = 12.89